OpenStudy (anonymous):

Given the function g(x) = −4x + 5, compare and contrast g(2) and g(−4). Choose the statement that is true concerning these two values. A. The value of g(2) is larger than the value of g(−4). B. The value of g(2) is smaller than the value of g(−4). C. The value of g(2) is the same as the value of g(−4). D. The values of g(2) and g(−4) cannot be compared.

OpenStudy (btaylor):

g(2) = -4(2) + 5 = -3 g(4) = -4(-4) + 5 = 21 so The value of g(2) is smaller than the value of g(−4). *the answer is (b)